# Marginally interpretable spatial logistic regression with bridge processes

### v2

@@ Acknowledgement
This research was partially supported by the National Institutes of Health (grant ID R01ES035625), by the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 856506), by the National Science Foundation (NSF IIS-2426762), and by the Office of Naval Research (N00014-21-1-2510). The authors thank Amy Herring and Georgia Papadogeorgou for helpful discussions.
@@ A.1 Proofs
Let $u=(u_{1},\dots,u_{n})$ be a finite realization from the bridge process where $R$ is a matrix with $(i,j)$th element $\mathcal{R}(s_{i},s_{j})$. By the law of total covariance, for any $i,j\in\{1,\dots,n\}$,
@@
$p\{Y_{ij},Z_{ij}\mid x_{ij},u(s_{i})\}$ =$=\{1(Z_{ij}>0)1(Y_{ij}=1)+1(Z_{ij}\leq 0)1(Y_{ij}=0)\}\sigma^{\prime}(\epsilon_{ij})$
@@ A.2 Details of posterior inference algorithm
We first introduce the notation. Write $N=\sum_{i=1}^{n}N_{i}$, let $X_{i}=[x_{i1}^{\mathrm{\scriptscriptstyle T}},\dots,x_{iN_{i}}^{\mathrm{\scriptscriptstyle T}}]^{\mathrm{\scriptscriptstyle T}}$ be a $N_{i}\times p$ matrix of predictors at the $i$th location, and $X=[X_{1}^{\mathrm{\scriptscriptstyle T}},\dots,X_{n}^{\mathrm{\scriptscriptstyle T}}]^{\mathrm{\scriptscriptstyle T}}$ be a $N\times p$ fixed effects design matrix. Also, let $Z=\mathrm{blockdiag}(1_{N_{1}},\dots,1_{N_{n}})$ be a $N\times n$ design matrix for random effects, so that the linear predictor vector becomes $X\beta+Zu$. Similarly, let $y_{i}=(y_{i1},\dots,y_{iN_{i}})$ be binary responses at location $i$ and $y=(y_{1}^{\mathrm{\scriptscriptstyle T}},\dots,y_{n}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ be a response vector of length $N$. Let $\Omega_{i}=\hbox{diag}(\omega_{i1},\dots,\omega_{iN_{i}})$ be a $N_{i}\times N_{i}$ diagonal matrix, $\Omega=\mathrm{blockdiag}(\Omega_{1},\dots,\Omega_{n})$ be a $N\times N$ diagonal matrix, and $\Omega_{nn}=Z^{\mathrm{\scriptscriptstyle T}}\Omega Z$ which is an $n\times n$ diagonal matrix with elements $\sum_{j=1}^{N_{i}}\omega_{ij}$ for $i=1,\dots,n$.
@@ A.2.1 Step 1
Conditional on $\Omega$, the likelihood is proportional to $\mathrm{N}_{N}\{\Omega^{-1}(y-0.51_{N});X\beta+Zu_{1:n},\Omega^{-1}\}$. Thus, integrating out $u_{1:n}\sim\mathrm{N}_{n}(0_{n},\lambda R)$, we have
$p(\beta\mid\omega,\lambda)\propto\mathrm{N}_{N}\{\Omega^{-1}(y-0.51_{N});X\beta,\Omega^{-1}+\lambda ZRZ^{\mathrm{\scriptscriptstyle T}}\}\times N_{p}(\beta;\mu_{\beta},\Sigma_{\beta}),$
which yields $(\beta\mid\omega,\lambda)\sim\mathrm{N}_{p}(Q_{1}^{-1}b_{1},Q_{1}^{-1})$ with $Q_{1}=X^{\mathrm{\scriptscriptstyle T}}(\Omega^{-1}+\lambda ZRZ^{\mathrm{\scriptscriptstyle T}})^{-1}X+\Sigma_{\beta}^{-1}$ and $b_{1}=\Sigma_{\beta}^{-1}\mu_{\beta}+X^{\mathrm{\scriptscriptstyle T}}(\Omega^{-1}+\lambda ZRZ^{\mathrm{\scriptscriptstyle T}})^{-1}\Omega^{-1}(y-0.51_{N})$. The inversion of $N\times N$ matrix $(\Omega^{-1}+\lambda ZRZ^{\mathrm{\scriptscriptstyle T}})$ can be done efficiently using the Woodbury formula
$(\Omega^{-1}+\lambda ZRZ^{\mathrm{\scriptscriptstyle T}})^{-1}=\Omega-\Omega Z(\lambda^{-1}R^{-1}+\Omega_{nn})^{-1}Z^{\mathrm{\scriptscriptstyle T}}\Omega$ (A.1)
@@
Suppose an independent normal scale mixture prior is used for $\beta$, say independent $t$ priors with degrees of freedom $\nu$, mean $\mu_{\beta}$ and scale $\sigma_{\beta}$. This corresponds to $\beta_{k}\mid\gamma_{k}\sim\mathrm{N}(0,\gamma_{k})$ and $1/\gamma_{k}\sim\mathrm{Ga}(\nu/2,\nu\sigma_{\beta}^{2}/2)$, independently for $k=1,\dots,p$, and one can add an additional sampling step $1/\gamma_{j}\sim\mathrm{Ga}(\nu/2+1/2,\nu\sigma_{\beta}^{2}/2+\beta^{2}/2)$ in the Algorithm.
@@ A.2.2 Step 2
We first derive Step 2(ii), the full conditional distribution of $u_{1:n}$. Given $\Omega$ and $\beta$, the likelihood $\mathrm{N}_{N}\left\{Zu_{1:n};\Omega^{-1}(y-0.51_{N})-X\beta,\Omega^{-1}\right\}$ can be recognized as a normal density in terms of $u_{1:n}$ since $Z$ has full column rank. This yields the full conditional distribution
@@
$p(u_{1:n}\mid\lambda,\omega,\beta)\propto\mathrm{N}_{n}[u_{1:n};\Omega_{nn}^{-1}Z^{\mathrm{\scriptscriptstyle T}}\Omega\{\Omega^{-1}(y-0.51_{N})-X\beta\},\Omega_{nn}^{-1}]\times\mathrm{N}_{n}(u_{1:n};0,\lambda R)$ (A.2)
@@
so $(u_{1:n}\mid\lambda,\omega,\beta)\sim\mathrm{N}_{n}(Q_{2}^{-1}b_{2},Q_{2}^{-1})$ with $Q_{2}=\Omega_{nn}+\lambda^{-1}R^{-1}$ and $b_{2}=Z^{\mathrm{\scriptscriptstyle T}}\{(y-0.51_{N})-\Omega X\beta\}$. For Step 2(i), from the expression (A.2), we can obtain a collapsed likelihood with the $u_{1:n}$ marginalized out. Thus, the target distribution of $\lambda$ and $\phi$ is given as
@@
$p(\lambda,\phi\mid\omega,\beta)\propto\underbrace{\mathrm{N}_{n}[\Omega_{nn}^{-1}Z^{\mathrm{\scriptscriptstyle T}}\Omega\{\Omega^{-1}(y-0.51_{N})-X\beta\};0,\Omega_{nn}^{-1}+\lambda R]}_{\mathcal{L}(\lambda)}\times{p_{\textsc{m}}}(\lambda;\phi)\times p(\phi)$ (A.3)
@@
and when $\lambda$ is integrated out, we have $p(\phi\mid\omega,\beta)\propto\int\mathcal{L}(\lambda){p_{\textsc{m}}}(\lambda;\phi)\mathrm{d}\lambda\times p(\phi)$. To jointly draw $\lambda$ and $\phi$ from (A.3), we use a particle marginal Metropolis-Hastings sampler (Andrieu et al.,, 2010). Let $L$ be the number of particles and $\lambda^{(1)},\dots,\lambda^{(L)}$ be the current set of particles. Then Step 2(i) proceeds as: (A) Draw candidate $\phi^{\star}\sim q(\phi^{\star}\mid\phi)$ with some proposal distribution. (B) Draw new particles independently from $\lambda^{\star(1)},\dots,\lambda^{\star(L)}\sim{p_{\textsc{m}}}(\lambda;\phi^{\star})$. (C) Draw a candidate $\lambda^{\star}$ among new particles $\lambda^{\star(1)},\dots,\lambda^{\star(L)}$ with probability proportional to $\mathcal{L}(\lambda^{\star(l)})$, $l=1,\dots,L$. (D) Accept $(\phi^{\star},\lambda^{\star})$ and set of particles $\{\lambda^{\star(l)}\}_{l=1}^{L}$ with probability
$\min\left\{1,\frac{\{\sum_{l=1}^{L}\mathcal{L}(\lambda^{\star(l)})\}p(\phi^{\star})}{\{\sum_{l=1}^{L}\mathcal{L}(\lambda^{(l)})\}p(\phi)}\times\frac{q(\phi\mid\phi^{\star})}{q(\phi^{\star}\mid\phi)}\right\}$
otherwise keep $(\phi,\lambda)$ and current set of particles $\{\lambda^{(l)}\}_{l=1}^{L}$.
@@
In all data analysis examples where the spatial range parameter $\rho$ is sampled together with $\phi$, we set $L=20$ and use coordinatewise logit transform to map parameters $(\phi,\rho)\in(0,1)\times(0.001,0.3)$ to $\mathbb{R}^{2}$, and utilize Metropolis-Hastings proposal adaptation settings as in Haario et al., (2001) to determine proposal distribution $q(\phi^{\star},\rho^{\star}\mid\phi,\rho)$.

### v3

@@ Acknowledgement
This research was partially supported by the National Institutes of Health (grant ID R01ES035625), by the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 856506), by the National Science Foundation (NSF IIS-2426762), and by the Office of Naval Research (N00014-24-1-2626). The authors thank anonymous reviewers, Amy Herring, and Georgia Papadogeorgou for helpful suggestions and discussions.
@@ A.1 Proofs
Proof of Proposition 1.
Let $\eta_{i}=\mathbf{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}$. Denote the marginal mean be $E(Y_{ij}\mid\mathbf{x}_{i})=\int_{-\infty}^{\infty}\operatorname*{logit}^{-1}(\eta_{i}+u_{i})\mathrm{d}u_{i}=\operatorname*{logit}^{-1}(\phi\eta_{i})=p_{i}$. We first show $\phi=E(\operatorname*{var}(Y_{ij}\mid\mathbf{x}_{i},u_{i}))/\operatorname*{var}(Y_{ij}\mid\mathbf{x}_{i})$. We have $E(\operatorname*{var}(Y_{ij}\mid\mathbf{x}_{i},u_{i}))=E[\{\operatorname*{logit}^{-1}(\eta_{i}+u_{i})\{1-\operatorname*{logit}^{-1}(\eta_{i}+u_{i})\}]=p_{i}-\int_{-\infty}^{\infty}\frac{(e^{u_{i}+\eta_{i}})^{2}}{(1+e^{u_{i}+\eta_{i}})^{2}}{p_{\textsc{br}}}(u_{i};\phi)\mathrm{d}u_{i}$, and the second term is
$\int_{-\infty}^{\infty}\frac{(e^{u_{i}+\eta_{i}})^{2}}{(1+e^{u_{i}+\eta_{i}})^{2}}{p_{\textsc{br}}}(u_{i};\phi)\mathrm{d}u_{i}$ $=\int_{-\infty}^{\infty}\frac{e^{u_{i}+\eta_{i}}(1+e^{u_{i}+\eta_{i}})-e^{u_{i}+\eta_{i}}}{(1+e^{u_{i}+\eta_{i}})^{2}}{p_{\textsc{br}}}(u_{i};\phi)\mathrm{d}u_{i}$
$=\frac{e^{\phi\eta_{i}}}{1+e^{\phi\eta_{i}}}-\frac{\phi e^{\phi\eta_{i}}}{(1+e^{\phi\eta_{i}})^{2}}$
$=p_{i}-\phi p_{i}(1-p_{i})$
where in the second equality, we used
$\int_{-\infty}^{\infty}\frac{e^{u+\eta}}{1+e^{u+\eta}}{p_{\textsc{br}}}(u;\phi)\mathrm{d}u=\frac{e^{\phi\eta}}{1+e^{\phi\eta}},\quad\int_{-\infty}^{\infty}\frac{e^{u+\eta}}{(1+e^{u+\eta})^{2}}{p_{\textsc{br}}}(u;\phi)\mathrm{d}u=\frac{\phi e^{\phi\eta}}{(1+e^{\phi\eta})^{2}}$
which follows from Wang and Louis, (2003) equations (2.1) and (2.2). Therefore, $E(\operatorname*{var}(Y_{ij}\mid\mathbf{x}_{i},u_{i}))=\phi p_{i}(1-p_{i})=\phi\operatorname*{var}(Y_{ij}\mid\mathbf{x}_{i})$. Moreover,
$\textstyle{\operatorname*{var}\left(\sum_{j=1}^{n_{i}}Y_{ij}\mid\mathbf{x}_{i}\right)}$ $=\textstyle{\operatorname*{var}\left(E\left(\sum_{j=1}^{n_{i}}Y_{ij}\mid\mathbf{x}_{i},u_{i}\right)\right)+E\left(\operatorname*{var}\left(\sum_{j=1}^{n_{i}}Y_{ij}\mid\mathbf{x}_{i},u_{i}\right)\right)}$
$=\operatorname*{var}\left(n_{i}E\left(Y_{ij}\mid\mathbf{x}_{i},u_{i}\right)\right)+E\left(n_{i}\operatorname*{var}\left(Y_{ij}\mid\mathbf{x}_{i},u_{i}\right)\right)$
$=n_{i}^{2}\operatorname*{var}\left(E\left(Y_{ij}\mid\mathbf{x}_{i},u_{i}\right)\right)+n_{i}E\left(\operatorname*{var}\left(Y_{ij}\mid\mathbf{x}_{i},u_{i}\right)\right)$
$=n_{i}^{2}(1-\phi)\operatorname*{var}(Y_{ij}\mid\mathbf{x}_{i})+n_{i}\phi\operatorname*{var}(Y_{ij}\mid\mathbf{x}_{i})$
$=n_{i}\operatorname*{var}(Y_{ij}\mid\mathbf{x}_{i})\{1+(n_{i}-1)(1-\phi)\}.$
@@
Let $u=(u_{1},\dots,u_{n})$ be a finite realization from the bridge process where $\mathbf{R}$ is a matrix with $(i,j)$th element $\mathcal{R}(s_{i},s_{j})$. By the law of total covariance, for any $i,j\in\{1,\dots,n\}$,
@@
$p\{Y_{ij},Z_{ij}\mid x_{ij},u(s_{i})\}$ $=\{1(Z_{ij}>0)1(Y_{ij}=1)+1(Z_{ij}\leq 0)1(Y_{ij}=0)\}\sigma^{\prime}(\epsilon_{ij})$
@@ A.2 Details of posterior inference algorithm
We first introduce the notation. Write $N=\sum_{i=1}^{n}N_{i}$, let $\mathbf{X}_{i}=[\mathbf{x}_{i1}^{\mathrm{\scriptscriptstyle T}},\dots,\mathbf{x}_{iN_{i}}^{\mathrm{\scriptscriptstyle T}}]^{\mathrm{\scriptscriptstyle T}}$ be a $N_{i}\times p$ matrix of predictors at the $i$th location, and $\mathbf{X}=[\mathbf{X}_{1}^{\mathrm{\scriptscriptstyle T}},\dots,\mathbf{X}_{n}^{\mathrm{\scriptscriptstyle T}}]^{\mathrm{\scriptscriptstyle T}}$ be a $N\times p$ fixed effects design matrix. Also, let $\mathbf{Z}=\mathrm{blockdiag}(\bm{1}_{N_{1}},\dots,\bm{1}_{N_{n}})$ be a $N\times n$ design matrix for random effects, so that the linear predictor vector becomes $\mathbf{X}\bm{\beta}+\mathbf{Z}\mathbf{u}_{1:n}$. Similarly, let $\mathbf{y}_{i}=(y_{i1},\dots,y_{iN_{i}})$ be binary responses at location $i$ and $\mathbf{y}=(\mathbf{y}_{1}^{\mathrm{\scriptscriptstyle T}},\dots,\mathbf{y}_{n}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ be a response vector of length $N$. Let $\bm{\Omega}_{i}=\hbox{diag}(\omega_{i1},\dots,\omega_{iN_{i}})$ be a $N_{i}\times N_{i}$ diagonal matrix, $\bm{\Omega}=\mathrm{blockdiag}(\bm{\Omega}_{1},\dots,\bm{\Omega}_{n})$ be a $N\times N$ diagonal matrix, and $\bm{\Omega}_{nn}=\mathbf{Z}^{\mathrm{\scriptscriptstyle T}}\bm{\Omega}\mathbf{Z}$ which is an $n\times n$ diagonal matrix with elements $\sum_{j=1}^{N_{i}}\omega_{ij}$ for $i=1,\dots,n$. Let $\mathbf{R}_{\rho}=[\mathcal{R}(s_{i},s_{j})]_{i,j}$ be an $n\times n$ correlation matrix with parameter $\rho$.
The Algorithm 2 describes a one cycle of a partially collapsed Gibbs sampler for a fully Bayesian approach. We first describe the details of the algorithm for a fully Bayesian approach, and specialize it to the Algorithm for an empirical Bayes approach.
[1] Sample $\bm{\beta}\sim[\bm{\beta}\mid\bm{\omega}^{\mathrm{(old)}},\lambda^{\mathrm{(old)}},\rho^{\mathrm{(old)}}]$ from multivariate normal, where $u_{1:n}$ is collapsed out.
[2] Jointly sample $(\phi,\rho,\lambda,\mathbf{u}_{1:n})\sim[\phi,\lambda,\mathbf{u}_{1:n}\mid\bm{\omega}^{\mathrm{(old)}},\bm{\beta}]$ in two steps:
   [2(i)] Sample $(\phi,\rho,\lambda)\sim[\phi,\rho,\lambda\mid\bm{\omega}^{\mathrm{(old)}},\bm{\beta}]$ using particle marginal Metropolis–Hastings,
   [2(ii)] Sample $\mathbf{u}_{1:n}\sim[\mathbf{u}_{1:n}\mid\phi,\lambda,\bm{\omega}^{\mathrm{(old)}},\bm{\beta}]$ from multivariate normal.
[3] Sample $\omega_{ij}\sim[\omega_{ij}\mid\bm{\beta},\mathbf{u}_{1:n}]$ from Pólya-Gamma, independently for all $i,j$.
Algorithm 2 One cycle of a partially collapsed Gibbs sampler.
@@ A.2.1 Step 1
Conditional on $\bf\Omega$, the likelihood is proportional to $\mathrm{N}_{N}\{\bm{\Omega}^{-1}(\mathbf{y}-0.5\bm{1}_{N});\mathbf{X}\bm{\beta}+\mathbf{Z}\mathbf{u}_{1:n},\bm{\Omega}^{-1}\}$. Thus, integrating out $\mathbf{u}_{1:n}\sim\mathrm{N}_{n}(\bm{0}_{n},\lambda\mathbf{R}_{\rho})$, we have
$p(\bm{\beta}\mid\bm{\omega},\lambda,\rho)\propto\mathrm{N}_{N}\{\bm{\Omega}^{-1}(\mathbf{y}-0.5\bm{1}_{N});\mathbf{X}\bm{\beta},\bm{\Omega}^{-1}+\lambda\mathbf{Z}\mathbf{R}_{\rho}\mathbf{Z}^{\mathrm{\scriptscriptstyle T}}\}\times N_{p}(\bm{\beta};\bm{\mu}_{\beta},\bm{\Sigma}_{\beta}),$
which yields $(\bm{\beta}\mid\bm{\omega},\lambda,\rho)\sim\mathrm{N}_{p}(\mathbf{Q}_{1}^{-1}\mathbf{b}_{1},\mathbf{Q}_{1}^{-1})$ with $\mathbf{Q}_{1}=\mathbf{X}^{\mathrm{\scriptscriptstyle T}}(\bm{\Omega}^{-1}+\lambda\mathbf{Z}\mathbf{R}_{\rho}\mathbf{Z}^{\mathrm{\scriptscriptstyle T}})^{-1}\mathbf{X}+\bm{\Sigma}_{\beta}^{-1}$ and $\mathbf{b}_{1}=\bm{\Sigma}_{\beta}^{-1}\bm{\mu}_{\beta}+\mathbf{X}^{\mathrm{\scriptscriptstyle T}}(\bm{\Omega}^{-1}+\lambda\mathbf{Z}\mathbf{R}_{\rho}\mathbf{Z}^{\mathrm{\scriptscriptstyle T}})^{-1}\bm{\Omega}^{-1}(\mathbf{y}-0.5\bm{1}_{N})$. The inversion of $N\times N$ matrix $(\bm{\Omega}^{-1}+\lambda\mathbf{Z}\mathbf{R}_{\rho}\mathbf{Z}^{\mathrm{\scriptscriptstyle T}})$ can be done efficiently using the Woodbury formula
$(\bm{\Omega}^{-1}+\lambda\mathbf{Z}\mathbf{R}_{\rho}\mathbf{Z}^{\mathrm{\scriptscriptstyle T}})^{-1}=\bm{\Omega}-\bm{\Omega}\mathbf{Z}(\lambda^{-1}\mathbf{R}_{\rho}^{-1}+\bm{\Omega}_{nn})^{-1}\mathbf{Z}^{\mathrm{\scriptscriptstyle T}}\bm{\Omega}$ (A.1)
@@
Suppose an independent normal scale mixture prior is used for $\bm{\beta}$, say independent $t$ priors with degrees of freedom $\nu$, mean $\mu_{\beta}$ and scale $\sigma_{\beta}$. This corresponds to $\beta_{k}\mid\gamma_{k}\sim\mathrm{N}(0,\gamma_{k})$ and $1/\gamma_{k}\sim\mathrm{Ga}(\nu/2,\nu\sigma_{\beta}^{2}/2)$, independently for $k=1,\dots,p$, and one can add an additional sampling step $1/\gamma_{k}\sim\mathrm{Ga}(\nu/2+1/2,\nu\sigma_{\beta}^{2}/2+\beta_{k}^{2}/2)$ in the Algorithm.
@@ A.2.2 Step 2
We first derive Step 2(ii), the full conditional distribution of $\mathbf{u}_{1:n}$. Given $\bm{\Omega}$ and $\bm{\beta}$, the likelihood $\mathrm{N}_{N}\left\{\mathbf{Z}\mathbf{u}_{1:n};\bm{\Omega}^{-1}(\mathbf{y}-0.5\bm{1}_{N})-\mathbf{X}\bm{\beta},\bm{\Omega}^{-1}\right\}$ can be recognized as a normal density in terms of $\mathbf{u}_{1:n}$ since $Z$ has full column rank. This yields the full conditional distribution
@@
$p(\mathbf{u}_{1:n}\mid\lambda,\omega,\bm{\beta},\rho)\propto\mathrm{N}_{n}[\mathbf{u}_{1:n};\bm{\Omega}_{nn}^{-1}\mathbf{Z}^{\mathrm{\scriptscriptstyle T}}\bm{\Omega}\{\bm{\Omega}^{-1}(\mathbf{y}-0.5\bm{1}_{N})-\mathbf{X}\bm{\beta}\},\bm{\Omega}_{nn}^{-1}]\times\mathrm{N}_{n}(\mathbf{u}_{1:n};\bm{0},\lambda\mathbf{R}_{\rho})$ (A.2)
@@
so $(\mathbf{u}_{1:n}\mid\lambda,\omega,\bm{\beta},\rho)\sim\mathrm{N}_{n}(\mathbf{Q}_{2}^{-1}\mathbf{b}_{2},\mathbf{Q}_{2}^{-1})$ with $\mathbf{Q}_{2}=\bm{\Omega}_{nn}+\lambda^{-1}\mathbf{R}_{\rho}^{-1}$ and $\mathbf{b}_{2}=\mathbf{Z}^{\mathrm{\scriptscriptstyle T}}\{(\mathbf{y}-0.5\bm{1}_{N})-\bm{\Omega}\mathbf{X}\bm{\beta}\}$. For Step 2(i), from the expression (A.2), we can obtain a collapsed likelihood with the $\mathbf{u}_{1:n}$ marginalized out. Thus, the target distribution of $\lambda$, $\phi$, and $\rho$ is given as
@@
$p(\lambda,\phi,\rho\mid\bm{\omega},\bm{\beta})\propto\underbrace{\mathrm{N}_{n}[\bm{\Omega}_{nn}^{-1}Z^{\mathrm{\scriptscriptstyle T}}\bm{\Omega}\{\bm{\Omega}^{-1}(\mathbf{y}-0.5\bm{1}_{N})-\mathbf{X}\bm{\beta}\};\bm{0},\bm{\Omega}_{nn}^{-1}+\lambda\mathbf{R}_{\rho}]}_{\mathcal{L}(\lambda,\rho)}{p_{\textsc{m}}}(\lambda;\phi)p(\phi)p(\rho)$ (A.3)
@@
and when $\lambda$ is integrated out, we have $p(\phi,\rho\mid\bm{\omega},\bm{\beta})\propto\int\mathcal{L}(\lambda,\rho){p_{\textsc{m}}}(\lambda;\phi)\mathrm{d}\lambda\times p(\phi)p(\rho)$. To jointly draw $\lambda$, $\rho$ and $\phi$ from (A.3), we use a particle marginal Metropolis-Hastings sampler (Andrieu et al., , 2010). Let $L$ be the number of particles and $\lambda^{(1)},\dots,\lambda^{(L)}$ be the current set of particles. Then Step 2(i) proceeds as: (A) Draw candidate $(\phi^{\star},\rho^{\star})\sim q(\phi^{\star},\rho^{\star}\mid\phi,\rho)$ with some proposal distribution. (B) Draw new particles independently from $\lambda^{\star(1)},\dots,\lambda^{\star(L)}\sim{p_{\textsc{m}}}(\lambda;\phi^{\star})$. (C) Draw a candidate $\lambda^{\star}$ among new particles $\lambda^{\star(1)},\dots,\lambda^{\star(L)}$ with probability proportional to $\mathcal{L}(\lambda^{\star(l)},\rho^{\star})$, $l=1,\dots,L$. (D) Accept $(\phi^{\star},\rho^{\star},\lambda^{\star})$ and set of particles $\{\lambda^{\star(l)}\}_{l=1}^{L}$ with probability
$\min\left\{1,\frac{\{\sum_{l=1}^{L}\mathcal{L}(\lambda^{\star(l)},\rho^{\star})\}p(\phi^{\star})p(\rho^{\star})}{\{\sum_{l=1}^{L}\mathcal{L}(\lambda^{(l)},\rho^{\star})\}p(\phi)p(\rho)}\times\frac{q(\phi,\rho\mid\phi^{\star},\rho^{\star})}{q(\phi^{\star},\rho^{\star}\mid\phi,\rho)}\right\}$
otherwise keep $(\phi,\rho,\lambda)$ and current set of particles $\{\lambda^{(l)}\}_{l=1}^{L}$.
@@
In fully Bayesian data analysis examples, we set $L=20$ and use coordinatewise logit transform to map parameters $(\phi,\rho)\in(0,1)\times(0.001,0.3)$ to $\mathbb{R}^{2}$, and utilize Metropolis-Hastings proposal adaptation settings as in Haario et al., (2001) to determine proposal distribution $q(\phi^{\star},\rho^{\star}\mid\phi,\rho)$.
Remark 1.
For the joint sampling of $(\lambda,\phi,\rho)$, the particle marginal Metropolis-Hastings sampler (Andrieu et al., , 2010) is suitable for this setting for the following reasons. First, the parameter $\phi$ only depends on $\lambda$, and conditional sampling between $\phi|\lambda,-]$ and $\lambda|\phi,-]$ leads to very high autocorrelation of $\phi$ and $\lambda$, and joint sampling of $\phi$ and $\lambda$ is highly desired. Next, from the joint density $p(\phi){p_{\textsc{m}}}(\lambda;\phi)$, the mixing density ${p_{\textsc{m}}}(\lambda;\phi)$ is represented as a alternating series (see Proposition A.2) and may unstable to evaluate, but it is easy to sample $\lambda$ using Theorem 1 and thus the joint density evaluation can be replaced with Monte Carlo estimates with particles, which is unbiased. Finally, using a simple Metropolis-Hastings proposal may suffer from a low acceptance probability, where the particle marginal Metropolis-Hastings algorithm leverages multiple proposals and improves the acceptance rate.